# Parameter Symmetry Breaking and Restoration Determines the Hierarchical Learning in AI Systems

### v1

@@ 5 Representation Learning Requires Parameter Symmetry
Representation learning is believed to be the most essential aspect of deep learning (Bengio et al., 2013). Learned representations of neural networks are found to take almost universally hierarchical forms, where earlier layers encode a large variety of low-level features and later layers learn a composed and abstract representation that is invariant to the changes in the low-level details (Zeiler & Fergus, 2014). One reason why symmetry may serve as a driving mechanism for learning these structured representations is that these structures almost always involve compressing information onto a few neurons and come with a low-rank structure in the hidden layer (Alain, 2016; Masarczyk et al., 2024; Xu et al., 2023; Papyan et al., 2020), and a primary low-rank mechanism for deep neural networks is through the permutation symmetry of the layer or the rescaling symmetry of ReLU neurons (Fukumizu, 1996; Fukumizu & Amari, 2000; Ziyin, 2024). As a primary example, neural collapses (NC) often happen in an image classification task, where the inner-class variations are found to disappear in the last and intermediate layers of the neural network (Papyan et al., 2020), resulting in a representation whose rank matches the number of classes. This implies that the network has learned a hierarchical representation, where, in the later layers, only high-level features are encoded. These results motivate the following hypothesis:
@@
Figure 4: Neural collapse (NC) only happens when permutation symmetry is present. NC is a primary example of how invariant high-level representations emerge in neural networks (Papyan et al., 2020) and exist quite generally in classification, regression, and large language models (Andriopoulos et al., 2024; Wu & Papyan, 2024). When NC happens, the learned representation must be low-rank; however, Ziyin et al. (2025b) showed that if the permutation symmetries are removed, the learned representation is always full-rank. This implies that permutation symmetry is a necessary condition for NC to happen. The figures show the representation alignment of $100$ CIFAR10 images across $10$ classes ($10$ images in each class). The color represents the correlation between representations. Left: The vanilla model exhibits neural collapse, where all neurons are similar for the same class. Right: The innerclass variation becomes significant when the permutation symmetry is removed.
@@
See Figure 4. Here, we train a standard ResNet18 on CIFAR-10, which is known to exhibit NC. In comparison, we also train a ResNet18 whose symmetries have all been removed using the method proposed in (Ziyin et al., 2025b). We see that after removing permutation symmetries, the innerclass variation of representations no longer vanishes. In Section A.4, we show that the spectral gap between class means and innerclass variations also becomes smaller.
More broadly, it is important to understand how representations build up as the input data passes through layers of a neural network. It seems likely that there are at least three distinctive regimes in the layers of a trained net: the first few layers of neural networks serve as an expansion phase where the representation becomes linearly separable (Alain, 2016), which requires the layer to be wide and implies a high rank (Nguyen et al., 2018); then, a “reduction” phase happens where the irrelevant information is thrown away and the neurons encode more and more compact information (Xu et al., 2023; Rangamani et al., 2023); lastly, a “transmission” phase where the layers do nothing except transmitting the signal it receives (Masarczyk et al., 2024). The NC can be seen as an example of the transmission and the simplest case of such a hierarchical representation (Ziyin et al., 2025a). These results all suggest that the representation rank is a good metric of the complexity of the layer. As Figure 5 shows, these changes in the representation ranks match well the symmetry-breaking level of the layer, a strong evidence that symmetry may drive the formation of hierarchical representations.
Figure 5: Neural networks learn a hierarchical representation. Recent works on representation learning have suggested that the rank of the latent representation first increases and then decreases through the layers (Xu et al., 2023; Masarczyk et al., 2024). This is reasonable because, on the one hand, a network needs to be wide enough to learn disconnected decision regions (Nguyen et al., 2018), while permutation symmetries drive towards low-rankness (Ziyin, 2024). This figure shows the rank and degree of symmetry breaking, which can be seen as the simplest metrics of the representation complexity, of different layers in a 5-layer FCN trained on CIFAR-10. This experiment shows hierarchical representations may be due to symmetry changes: the beginning layers feature symmetry breaking, and later layers are primarily symmetry restoration.
@@
The neural collapse is also a special case of what we will call “universal representation learning.” Recent works found that the representations of learned models are found to be universally aligned to different models trained on similar datasets (Bansal et al., 2021; Kornblith et al., 2019), and to the biological brains (Yamins et al., 2014). This interesting phenomenon has a rather philosophical undertone and has been termed “Platonic Representation Hypothesis” (Huh et al., 2024). Here, we say that the two neural networks have learned a universal representation if for all $x_{1},\ x_{2}$,
@@
Recent results have suggested that symmetry may play a key role in shaping the representation of neural networks. The following theorem can be proved by leveraging Theorem 5.4 of Ziyin et al. (2024), which shows that the double rotation symmetry in deep linear networks leads to the emergence of universal representations in different models, even when data and models are different.
@@
(Informal) Let $\mathcal{D}_{M}=\{(Mx_{i},y_{i})\}_{i}$ be the training set, where $M$ is an invertible matrix. Let deep linear networks A and B have arbitrary depths and widths. Let model A train with SGD on $\mathcal{D}_{M}$ and model B with SGD on $\mathcal{D}_{M^{\prime}}$ for arbitrary $M$, $M^{\prime}$. Then, if the model converges and is at the global minimum, every layer $h_{A}$ of A is perfectly aligned with every layer $h_{B}$ of B for any $x$:
$h_{A}(x)\propto Rh_{B}(x),\vspace{-2mm}$ (6)
@@
See Figure 6. This is an extraordinary fact: due to the double rotation symmetry, there exist infinitely many global minima for a deep linear network such that the representations are not aligned. Yet, SGD training finds a special solution such that every layer of $A$ is aligned with every layer of $B$ despite having different initializations, different data transformations (controlled by the arbitrary matrix $M$), and different architectures (width and depth). This is only possible if the first layer transforms the representation into an input-independent form. See Figure 6 for a demonstration. This theorem is a direct (perhaps the first) proof of the platonic representation hypothesis, implying that for any $x_{1},x_{2}$, Eq. (5) holds. Importantly, the mechanism does not belong to any previously conjectured mechanisms (capacity, simplicity, multitasking (Huh et al., 2024)). This example has nothing to do with multitasking. The result holds for any deep linear network, all having the same capacity and the same level of simplicity because all solutions parametrize the same input-output map. Here, the cause of the universal representation is symmetry alone: in the degenerate manifold of solutions, the training algorithm prefers a particular and universal one. This example showcases how symmetry is indeed an overlooked fundamental mechanism in deep learning.
An important future step is to generalize it to nonlinear networks. Due to extensive connectivity among the global minima of overparametrized networks (Nguyen, 2019), there is a great potential that parameter symmetry (in exact or approximate forms) plays a similar role in overparameterized neural networks (Zhao et al., 2022, 2023a).
Figure 6: Universally aligned representations emerge in differently trained neural networks. Many recent works demonstrate that different trained neural networks learn representations that are similar to each other and even to the biological brain (Huh et al., 2024; Kornblith et al., 2019; Yamins et al., 2014). Parameter symmetry may be a core mechanism for this intriguing phenomenon as it implies a universal alignment between representations (Ziyin et al., 2024). The figure shows the representations of two deep linear networks independently trained on randomly transformed MNIST become perfectly aligned for every pair of layers. The figure shows the average alignment between the same or different layers of two networks. This alignment does not weaken even if the input is arbitrarily transformed (Theorem 2). The black dashed line shows the average alignment to the input data, which is significantly weaker.
@@ 6 Mechanism and Control
A primary known mechanism is regularization in explicit or implicit forms. Since symmetry breaking is just the lack of symmetry restoration, we may focus on symmetry restoration. In explicit form, a simple weight decay has been shown to universally turn symmetric solutions energetically favorable to symmetry-broken solutions (Ziyin, 2024), in a manner similar to phase transition in physics (Landau & Lifshitz, 2013). In implicit form, the stochasticity in SGD is known to lead to an “implicit” regularization effect, sometimes similar to that of weight decay (Kunin et al., 2021; Chen et al., 2023). Another mechanism is data augmentation, as it can also be seen as a form of regularization (Dao et al., 2019). While understanding these mechanisms is an open problem, the easiest way to study it is to consider the effective loss landscape (also known as the “modified loss”) under SGD training (Geiping et al., 2021; Smith et al., 2021):
@@
where $L_{0}$ has the symmetry under consideration, $\Sigma$ is the gradient covariance matrix due to SGD sampling, and $T=\eta/S$ is the learning-rate-to-batch-size ratio. The data augmentation term can also be controlled by, say, a variance term $\sigma$. A careful inspection suggests that none of these terms breaks the $Z_{2}$ symmetries of $L_{0}$. So all $\sigma,\ \gamma,\ T$ can serve as regularizers that can induce a “spontaneous symmetry breaking,” which means that the symmetric states are energetically favored solutions at a critical regularization strength (Ziyin, 2024). We experimentally demonstrate these effects in Section A.3.
With explicit regularization, whether symmetry breaking or restoration happens depends on the competition between regularization $\gamma$ and the data signal. The more curious situation is when $\gamma=0$ and $T>0$. This type of symmetry breaking is more like a thermodynamic phase transition in physics as it is induced by an entropic force due to the SGD noise. Why do noisy dynamics favor symmetric states? One intuition is that the “temperature” of the dynamics is parameter-dependent, and stochastic dynamics tend to move to places that are “cold,” a common phenomenon in nature (Duhr & Braun, 2006; Anzini et al., 2019).  [^1] So, implicit regularization effects can be viewed as a thermal current obeying laws of nonequilibrium physics (Seifert, 2008). In other words, temperature gradient tends to create a flow. In the case of SGD dynamics on a neural network, the symmetric solutions are “ colder” because they have a lower-rank $\Sigma$ than the symmetry-broken solutions. This discussion offers a likely intuition for the mechanism, but a more formal treatment of the argument is still lacking. Different training techniques likely favor different types of symmetries, and one may leverage symmetry to understand the inductive biases of deep learning techniques.
@@
If hierarchies and complexity adaptivity are desired, one can introduce artificial symmetries to the neural network. A straightforward way to introduce symmetries is by introducing an additional parameter $v$ and multiplying it with the parameters $W$ for which a hierarchy is desired. For example, changing a layer weight from $W$ to $vW$ introduces a new symmetry without affecting the model expressivity: $|v|=\|W\|=0$ is the symmetric state, and the model will naturally adapt to break or restore this symmetry. This is a special case of a class of algorithms that introduce artificial symmetries to constrain the model parameters. (for example, see the DCS algorithm in Ziyin (2024); also, see Kolb et al. (2025)). A simple way to remove symmetries is to add some static randomness to the model, which has been found to help with the model connectivity (Lim et al., 2024) and avoid saddle points (Ziyin et al., 2025b). Finding the right symmetries to introduce or remove can be a fruitful future direction.
@@ 7 Outlook and Alternative Views
In this work, we have argued and demonstrated that a wide range of seemingly unrelated hierarchies emergent in AI systems is actually related to, if not directly caused by, the symmetry of the trained models. Symmetry breaking and restoration are fundamental mechanisms in physics relevant to the dynamics of almost every scale of nature – from scattering of the quarks to the large-scale structure formation of the universe (Miller et al., 1990; Preskill et al., 1991). If different symmetries govern the universal laws at different scales of nature (Anderson, 1972), it is natural to hypothesize that it may also lead to universal mechanisms and laws of learning in both artificial and biological systems. We suggest three primary hypothetical mechanisms that govern three fundamental aspects of hierarchical learning of neural networks, and validating or falsifying their related hypotheses is the most important next step. From a broader perspective, that symmetry can play such important roles in learning also calls for more interdisciplinary study of intelligence from the perspective of physics.
@@
In most of our examples, it is a combination of symmetry and some other effect that leads to the phenomenon. For example, the compression bias of SGD is a result of symmetry and training noise. Neural collapse is due to symmetry and regularization (Rangamani & Banburski-Fahey, 2022). Therefore, it is possibly the case that symmetry plus some form of explicit (weight decay) or implicit (noise, GD training, data augmentation) regularization is the dominating factor, and this may be an important direction of future research.
@@
Behind symmetry there may be more:
There are also possible more fundamental mechanisms than symmetry, which can explain what happens with symmetry and what happens without it. One such example could be the topology of the system. In fact, in theoretical physics, symmetry is often regarded as a special case of topology (for example, the rank of the Fisher information can be a good the topological invariant). This makes it possible to generalize our arguments to a wider range of systems. Some works have explored the role of topology in AI, but often in a scenario where symmetry is also present (Barannikov et al., 2020; Nurisso et al., 2024). Another interesting future direction is understanding what topology may achieve without requiring symmetries.
Acknowledgment
ILC acknowledges support in part from the Institute for Artificial Intelligence and Fundamental Interactions (IAIFI) through NSF Grant No. PHY-2019786. This work was also supported by the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF - 1231216.
@@ A.2 Learning Dynamics
In Figure 2, we use a teacher-student setting, where both the teacher and student are five-layer fully connected networks (FCNs) with 64 units per layer and tanh activation. The input and output dimensions are $10$ and $1$, respectively. Teacher network weights follow Kaiming initialization, and the output is scaled by 10. The student network is initialized near zero. The training dataset consists of $300$ samples generated as $\sin(\mathcal{N}(0,\sigma^{2}))$, with $\sigma$ ranging from 00 to $3$. The test dataset contains 500 samples from the same distribution. We train the student network with gradient descent (no weight decay) for $50,000$ iterations using the mean squared error (MSE) loss.
To calculate the symmetry-breaking distance, we sort the first-layer neurons by their norms and compute the $L_{2}$ distance between consecutive neurons. Figure 2 shows distances for $63$ neuron pairs in the first layer (see Section A.1). Results for quadratic and sin activations are presented in Figures 7 and 8.

### v2

@@ 5 Representation Learning Requires Parameter Symmetry
Figure 4: Neural networks learn a hierarchical representation. Recent works on representation learning have suggested that the rank of the latent representation first increases and then decreases through the layers [72, 47]. This is reasonable because, on the one hand, a network needs to be wide enough to learn disconnected decision regions [53], while permutation symmetries drive towards low-rankness [83]. This figure repeats the experiment in Ref. [47] and shows the rank and degree of symmetry breaking, which can be seen as the simplest metrics of the representation complexity, of different layers in a 5-layer FCN trained on CIFAR-10. This experiment shows hierarchical representations may be due to symmetry changes: the beginning layers feature symmetry breaking, and later layers are primarily symmetry restoration.
Representation learning is believed to be the most essential aspect of deep learning [10]. Learned representations of neural networks are found to take almost universally hierarchical forms, where earlier layers encode a large variety of low-level features and later layers learn a composed and abstract representation that is invariant to the changes in the low-level details [77]. One reason why symmetry may serve as a driving mechanism for learning these structured representations is that these structures almost always involve compressing information onto a few neurons and come with a low-rank structure in the hidden layer [2, 47, 72, 54], and a primary low-rank mechanism for deep neural networks is through the permutation symmetry of the layer or the rescaling symmetry of ReLU neurons [23, 24, 83]. As a primary example, neural collapses (NC) often happen in an image classification task, where the inner-class variations are found to disappear in the last and intermediate layers of the neural network [54], resulting in a representation whose rank matches the number of classes. This implies that the network has learned a hierarchical representation, where, in the later layers, only high-level features are encoded. These results motivate the following hypothesis:
@@
Figure 5: Neural collapse (NC) only happens when permutation symmetry is present. NC is a primary example of how invariant high-level representations emerge in neural networks [54] and exist quite generally in classification, regression, and large language models [5, 71]. When NC happens, the learned representation must be low-rank; however, Ref. [88] showed that if the permutation symmetries are removed, the learned representation is always full-rank. This implies that permutation symmetry is a necessary condition for NC to happen. The figures show the representation alignment of $100$ CIFAR10 images across $10$ classes ($10$ images in each class) and illustrates the result of Ref.  [88]. The color represents the correlation between representations. Left: The vanilla model exhibits neural collapse, where all neurons are similar for the same class. Right: The innerclass variation becomes significant when the permutation symmetry is removed.
@@
See Figure 5. Here, we train a standard ResNet18 on CIFAR-10, which is known to exhibit NC. In comparison, we also train a ResNet18 whose symmetries have all been removed using the method proposed in Ref. [88]. We see that after removing permutation symmetries, the innerclass variation of representations no longer vanishes. In Section A.4, we show that the spectral gap between class means and innerclass variations also becomes smaller.
More broadly, it is important to understand how representations build up as the input data passes through layers of a neural network. It seems likely that there are at least three distinctive regimes in the layers of a trained net: the first few layers of neural networks serve as an expansion phase where the representation becomes linearly separable [2], which requires the layer to be wide and implies a high rank [53]; then, a “reduction” phase happens where the irrelevant information is thrown away and the neurons encode more and more compact information [72, 59]; lastly, a “transmission” phase where the layers do nothing except transmitting the signal it receives [47]. The NC can be seen as an example of the transmission and the simplest case of such a hierarchical representation [84]. These results all suggest that the representation rank is a good metric of the complexity of the layer. As Figure 4 shows, these changes in the representation ranks match well the symmetry-breaking level of the layer, a strong evidence that symmetry may drive the formation of hierarchical representations.
@@
The neural collapse is also a special case of what we will call “universal representation learning.” Recent works found that the representations of learned models are found to be universally aligned to different models trained on similar datasets [8, 36], and to the biological brains [75]. This intriguing phenomenon has a philosophical undertone and has been termed “Platonic Representation Hypothesis” [30]. Here, we say that the two neural networks have learned a universal representation if for all $x_{1},\ x_{2}$,
@@
Recent results have suggested that symmetry may play a key role in shaping the representation of neural networks. The following theorem from Ref. [87] shows that the double rotation symmetry in deep linear networks leads to the emergence of universal representations in different models, even when data and models are different.
@@
(Informal, [87, Theorem 8]) Let $\mathcal{D}_{M}=\{(Mx_{i},y_{i})\}_{i}$ be the training set, where $M$ is an invertible matrix. Let deep linear networks A and B have arbitrary depths and widths. Let model A train with SGD on $\mathcal{D}_{M}$ and model B with SGD on $\mathcal{D}_{M^{\prime}}$ for arbitrary $M$, $M^{\prime}$. Then, if the model converges and is at the global minimum, every layer $h_{A}$ of A is perfectly aligned with every layer $h_{B}$ of B for any $x$:
$h_{A}(x)\propto Rh_{B}(x),$ (6)
@@
See Figure 6. This is an extraordinary fact: due to the double rotation symmetry, there exist infinitely many global minima for a deep linear network such that the representations are not aligned. Yet, SGD training finds a special solution such that every layer of $A$ is aligned with every layer of $B$ despite having different initializations, different data transformations (controlled by the arbitrary matrix $M$), and different architectures (width and depth). This is only possible if the first layer transforms the representation into an input-independent form. See Figure 6 for a demonstration. This theorem is a direct (perhaps the first) proof of the platonic representation hypothesis, implying that for any $x_{1},x_{2}$, Eq. (5) holds. Importantly, the mechanism does not belong to any previously conjectured mechanisms (capacity, simplicity, multitasking [30]). This example has nothing to do with multitasking. The result holds for any deep linear network, all having the same capacity and the same level of simplicity because all solutions parametrize the same input-output map. Here, the cause of the universal representation is symmetry alone: in the degenerate manifold of solutions, the training algorithm prefers a particular and universal one. This example showcases how symmetry is indeed an overlooked fundamental mechanism in deep learning.
Figure 6: Universally aligned representations emerge in differently trained neural networks. Many recent works demonstrate that different trained neural networks learn representations that are similar to each other and even to the biological brain [30, 36, 75]. Parameter symmetry may be a core mechanism for this intriguing phenomenon as it implies a universal alignment between representations [86]. The figure is reproduced from Ref. [87] and shows the representations of two deep linear networks independently trained on randomly transformed MNIST become perfectly aligned for every pair of layers. The figure shows the average alignment between the same or different layers of two networks. This alignment does not weaken even if the input is arbitrarily transformed (Theorem 2). The black dashed line shows the average alignment to the input data, which is significantly weaker.
An important future step is to generalize it to nonlinear networks. Due to extensive connectivity among the global minima of overparametrized networks [52], there is a great potential that parameter symmetry (in exact or approximate forms) plays a similar role in overparameterized neural networks [79, 80].
@@ 6 Mechanism and Control
Since symmetry breaking is just the lack of symmetry restoration, we may focus on restoration. A primary known mechanism for symmetry restoration is regularization in explicit or implicit forms. In explicit form, a simple weight decay has been shown to universally turn symmetric solutions energetically favorable to symmetry-broken solutions [83], in a manner similar to phase transition in physics [39]. In implicit form, the stochasticity in SGD is known to lead to an “implicit” regularization effect, sometimes similar to that of weight decay [38, 12]. Another mechanism is data augmentation, as it can also be seen as a form of regularization [15]. While understanding these mechanisms is an open problem, the easiest way to study it is to consider the entropic loss function (also known as the “modified loss”) under SGD training [27, 63, 87]:
@@
where $L_{0}$ has the symmetry under consideration, $\Sigma$ is the gradient covariance matrix due to SGD sampling, and $T=\eta/S$ is the learning-rate-to-batch-size ratio. The data augmentation term can also be controlled by, say, a variance term $\sigma$. A careful inspection suggests that none of these terms breaks the $Z_{2}$ symmetries of $L_{0}$. So all $\sigma,\ \gamma,\ T$ can serve as regularizers that can induce a “spontaneous symmetry breaking,” which means that the symmetric states are energetically favored solutions at a critical regularization strength [83]. We experimentally demonstrate these effects in Section A.3.
With explicit regularization, whether symmetry breaking or restoration happens depends on the competition between regularization $\gamma$ and the data signal. The more curious situation is when $\gamma=0$ and $T>0$. This type of symmetry breaking is more like a thermodynamic phase transition in physics as it is induced by an entropic force due to the SGD noise. Why do noisy dynamics favor symmetric states? One intuition is that the “temperature” of the dynamics is parameter-dependent, and stochastic dynamics tend to move to places that are “cold,” a common phenomenon in nature [19, 6]. In other words, temperature gradient tends to create a flow. In the case of SGD dynamics on a neural network, the symmetric solutions are “ colder” because they have a lower-rank $\Sigma$ than the symmetry-broken solutions. This discussion offers a likely intuition for the mechanism, but a more formal treatment of the argument is still lacking. Different training techniques likely favor different types of symmetries, and one may leverage symmetry to understand the inductive biases of deep learning techniques.
@@
If hierarchies and complexity adaptivity are desired, one can introduce artificial symmetries to the neural network. A straightforward way to introduce symmetries is by introducing an additional parameter $v$ and multiplying it with the parameters $W$ for which a hierarchy is desired. For example, changing a layer weight from $W$ to $vW$ introduces a new symmetry without affecting the model expressivity: $|v|=\|W\|=0$ is the symmetric state, and the model will naturally adapt to break or restore this symmetry. This is a special case of the DCS algorithm proposed in Ref. [83]. A simple way to remove symmetries is to add some static randomness to the model, which has been found to help with the model connectivity [44] and avoid saddle points [88]. Finding the right symmetries to introduce or remove can be a fruitful future direction.
@@ 7 Outlook and Alternative Views
In this position paper, we have leveraged existing results to argue that parameter symmetry is a key for unifying theories of deep learning because a wide range of seemingly unrelated hierarchies emergent in AI systems is actually related to, if not directly caused by, the symmetry of the trained models. Symmetry breaking and restoration are fundamental mechanisms in physics relevant to the dynamics of almost every scale of nature – from scattering of the quarks to the large-scale structure formation of the universe [48, 57]. If different symmetries govern the universal laws at different scales of nature [4], it is natural to hypothesize that it may also lead to universal mechanisms and laws of learning in both artificial and biological systems. We suggest three primary hypothetical mechanisms that govern three fundamental aspects of hierarchical learning of neural networks, and validating or falsifying their related hypotheses is the most important next step. From a broader perspective, that symmetry can play such important roles in learning also calls for more interdisciplinary study of intelligence from the perspective of physics.
@@
There is no unified explanation of deep learning:
This may very well be the case, and even laws of physics cannot explain every observation in nature. However, it is still highly possible that a small set of laws is sufficient to explain a large fraction of observations. For example, it is well possible that there may exist ten minimal laws with which we can explain 95% of the deep learning phenomenon, and if that is the case, symmetry may be one of those laws.
@@
In most of our examples, it is a combination of symmetry and some other effect that leads to the phenomenon. For example, the compression bias of SGD is a result of symmetry and training noise. Neural collapse is due to symmetry and regularization [58]. Therefore, it is possibly the case that symmetry plus some form of explicit (weight decay) or implicit (noise, GD training, data augmentation) regularization is the dominating factor, and this may be an important direction of future research.
@@ A.2 Learning Dynamics
In Figure 2, we use a teacher-student setting, where both the teacher and student are five-layer fully connected networks (FCNs) with 64 units per layer and tanh activation. The input and output dimensions are $10$ and $1$, respectively. Teacher network weights follow Kaiming initialization, and the output is scaled by 10. The student network is initialized with different scales. The training dataset consists of $300$ samples generated as $\sin(\mathcal{N}(0,\sigma^{2}))$, with $\sigma$ ranging from 00 to $3$. The test dataset contains 500 samples from the same distribution. We train the student network with gradient descent (no weight decay) for $50,000$ iterations using the mean squared error (MSE) loss.
To calculate the symmetry-breaking distance, we sort the first-layer neurons by their norms and compute the $L_{2}$ distance between consecutive neurons. Results for quadratic and sin activations are presented in Figures 7 and 8, including distances for $63$ neuron pairs in the first layer (see Section A.1).